# Epistemic Boundaries and Quantum Indeterminacy: What Local Observers Can (Not) Predict

### v2

@@
Epistemic Boundaries and Quantum Indeterminacy: What Local Observers Can (Not) Predict
Johannes Fankhauser johannes.j.fankhauser@gmail.com Institute for Theoretical Physics, University of Innsbruck, Faculty of Philosophy, University of Oxford
((Dated: December 23, 2023))
@@
One of quantum theory’s salient lessons is its inherent indeterminacy. That is, generic physical states imply uncertainty for the outcomes of measurements. We formally define and address whether quantum uncertainty could be fundamental or whether post-quantum theories can offer predictive advantage whilst conforming to the Born rule on average. We present a no-go claim combining three aspects: predictive advantage, signal-locality, and the epistemic relationship between quantum observers. The results of the analysis lead to the conclusion that there exists a fundamental limitation on genuine predictive advantage over standard quantum probabilities. However, we uncover a fascinating possibility: when the assumption of ‘reliable intersubjectivity’ between different observers is violated, subjective predictive advantage can, in principle, exist. This, in turn, entails an epistemic boundary between different observers of the same theory. The findings reconcile us to quantum uncertainty as an aspect of limits on Nature’s predictability.
Introduction
@@
Insofar as the formalism is deemed valid, quantum theory’s predictions are generally probabilistic. That is, generic physical states imply uncertainty for the outcomes of measurements. At face value, the quantum formalism suggests that its probabilism is unsurmountable, and so quantum indeterminacy has become one of the characteristics of the theory.
@@
More formally, quantum theory satisfies a ‘decorrelation principle’: Whenever a pure state is assigned to a target system $S$, the outcomes of any measurement are statistically independent of its environment. Therefore, the maximum information encoded in its environment is the information about a target system’s pure quantum state:
$p(x,y|\rho)=p(x|\rho_{S})\cdot p(y|\rho_{E}).$
@@
The question arises whether we expect quantum uncertainty to be fundamental in future theories as well, or whether their could exist non-quantum variables with predictive advantage.
@@
But why does the quantum formalism suggest that there is a limit to the predictability of an event? Is it because Nature is fundamentally indeterministic? Even if Nature were in fact deterministic, this raises the question why it shouldn’t be predictable. Or is quantum indeterminacy a consequence of a more fundamental physical principle? Or perhaps the question is ill-posed since the quantum state was never supposed to describe physical reality at all?
@@
In 1935 the puzzle about quantum theory’s incompleteness culminated in the notorious ‘EPR-argument’: Imagine a widely separated pair of an entangled particles that can thus no longer interact with one another. In turn, the situation is said to satisfy Einstein, Podolsky, and Rosen’s element of reality: Since “without in any way disturbing the system” (as the particles are space-like separated) “we can predict with certainty (i.e. with probability equal to unity) the value of a physical quantity, […] there exists an element of physical reality corresponding to this physical quantity.” [15, p. 777]. Since the particles are entangled, one can therefore simultaneously perform a position measurement on one and a momentum measurement on the other particle. If the remote measurement outcome can be predicted with certainty (due to the perfect quantum correlations) without interfering with the system, the outcome must be pre-determined and thus the quantum description is incomplete: Einstein was “[…] firmly convinced that the essentially statistical character of contemporary quantum theory is solely to be ascribed to the fact that this [theory] operates with an incomplete description of physical systems”. [15, p. 666]
@@
Einstein played with other gedanken experiments trying to conceive of scenarios violating Heisenberg’s uncertainty relations (see e.g. perpetua mobilia like the Einstein paradox, [7]). He was aware that knowledge of the uncertainty principles doesn’t imply that a more complete theory is impossible. Einstein was primarily concerned with the ontological rather than epistemic completeness of quantum mechanics. EPR never claimed that the theory would be predictively complete (see, for instance, [14]).
In his response to their argument Bohr sharply deviates from EPR’s suggestion that a more complete description of physical reality is possible. Not only could there not be more elements of reality than the wavefunction, but Bohr also denied that it wouldn’t describe reality at all. [7, p. 199-241]
Any appeal to the idea that nothing else than the quantum wave function exists akin to Bohr’s radical dismissal stops the quest for completeness right from the outset.
However, the question we are asking here is what can be inferred about predictability from the empirical predictions of quantum mechanics alone. That is, we seek out a possible argument for quantum completeness that does not stem from an a priori commitment to the wave function as the final means of making predictions.
One motivation is practical: What is the best possible prediction for physical phenomena that can be made? Is quantum mechanics (at least in principle) the most powerful prediction-making algorithm? What is the maximally achievable information of future measurement outcomes that can be encoded in the environment?
Secondly, future experiments may demonstrate that quantum theory isn’t applicable or valid in certain regimes. If that turns out to be true, it’s desirable to have physical principles that can be used to make statements about the predictability in post-quantum theories. That is, an understanding of predictability could be reached independently of the quantum framework, which may also inform future theory construction.
Thirdly, are there objective probabilities in nature? If yes, are those the quantum probabilities? That is, are quantum predictions irreducible?
The question of empirical completeness is interesting for two major reasons: A) It’s unclear—based on the empirical export of quantum theory alone—what all the variables in Nature are that determine a measurement outcome. And B) Can we gain a better understanding of quantum indeterminacy motivated by first principles?
We are thus not concerned with the ontological or metaphysical status of the quantum state, but rather with its epistemic role as a means for making predictions. Thus, the focus is shifted away from the ontological underpinning of the theory—such as e.g. in hidden variable theories—to prediction-making. [^1] Note that there exists a similar but distinct notion of completeness which is concerned with the question whether the quantum state is ontic. In other words, whether the same state of reality is compatible with more than one wavefunction. (see Pusey, Barrett, and Rudolph’s seminal result on the reality of the quantum state. [36]). This, in turn, supports the need for a clear distinction between what’s manifest, i.e. empirically accessible, from what isn’t, i.e. non-manifest. The variables that may determine a single measurement result don’t necessarily coincide with the ones utilisable for predicting that result. Rather than to attempt restoring determinism in quantum mechanics, the question is whether quantum mechanics is empirically complete.
In this paper we employ a strong correspondence between quantum non-locality and predictive completeness in an impossibility result taking into account the quantum reality of macroscopic observers. We show that quantum probabilities pose strong constraints on the intersubjectivity of information used to make predictions if observers are assumed to be local: The quantum state provides complete information about the predictions of measurement outcomes. That is, no predictive advantage can objectively exist for all observers of a theory reproducing the quantum probabilities on average. Moreover, predictive advantage is possible only when there exist ‘epistemic boundaries’ between observers.
Predictions of signal-local observers
@@
Let’s begin with an instructive example from classical physics. Imagine a classical thermodynamic system of a gas in a box to illustrate how empirical incompleteness may turn out. The box of gas may be described as a closed system with a constant temperature. Subject to kinetic gas theory, a measurement of the velocity $v$ of the gas particles in the ensemble is expected to give values distributed according to Maxwell-Boltzmann statistics $f_{MB}(v)$, peaked around some mean velocity. Thus, the velocity $v$ of an individual particle will be found with some likelihood $p(v)$. But let’s imagine the outcomes of many repeated measurements on the same gas yield doubly peaked distributions corresponding to two Maxwell-Boltzmann distributions of different mean velocities. And let’s assume the two peaks originate from the fact that there are two distinct gases in the box, namely nitrogen and oxygen. Since the two gases have different kinetic properties, this leads to separate mean velocities. But to an experimenter without a physical theory distinguishing the two molecules, only one gas is in the box. Thus, the experimenter’s original theory is empirically incomplete since the measurement outcomes will show an in-principle uncertainty resulting from treating the gas as one single substance. A more advanced theory that can distinguish nitrogen from oxygen will obtain more complete predictions. [^2] For this to be possible, an actual experiment, i.e. preparation, has to exist that distinguishes nitrogen from oxygen, such that, on average, the total distribution $f_{MB}(v)$ is still recovered. This is crucial, for otherwise, the improved theory wouldn’t be an empirical completion but rather a metaphysical completion.
@@
All variables involved—outcomes $x$, and preparations $y_{1},y_{2}$—are experimentally accessible variables on which the state preparations and predictions supervene. [^3] It is worth noting that indeed every probabilistic physical theory can be supplemented by a deterministic underpinning. (for any sequence of events there exists a descriptions that contains all information in the initial state of the system, cf. e.g. [18], [10, Chapter 3.4.1]. A viable version of determinism will thus involve restrictions on the domain of variables allowed and the complexity of laws invoked). For instance, quantum theory is explicitly probabilistic, but pilot wave theory provides a deterministic underpinning of it.
@@
The idea to connect (a version of) locality with quantum completeness dates back to Einstein, Podolsky and Rosen, who argued for the incompleteness of quantum theory based on the existence of entangled states and their non-local properties [15]. If quantum measurement outcomes aren’t further pre-determined by more than the probabilistic quantum prediction, they must be non-local to recover the perfect correlations of entangled states. Supplementing the quantum description with so-called hidden variables could then perhaps restore determinism, thus completing the theory to avoid what is often called ‘action at a distance’.
@@
But by Bell’s theorem, any such completions must be non-local too.[5, 33] By bringing in quantum non-locality in the sense of Bell the question of empirical completeness becomes in fact tractable: To avoid the paradoxical challenges of superluminal signalling in a so extended theory, the theory can either not be completed, or some mechanism exists limiting the extend to which such variables are empirically accessible. For otherwise if experimenters could control said variables, they could employ them to signal.
@@
The basic setup of the thought experiment used in the main result consists of three observers Alice, Bob, and Maggie which we shall abbreviate with $A,B$, and $M$. The source $S$ emits a pair of spin-$\frac{1}{2}$ particles prepared in the pure entangled state $\rho=\ket{\psi}\bra{\psi}$ | with |$\ket{\psi}:=\alpha\ket{00^{\prime}}+\sqrt{1-\alpha^{2}}\ket{11^{\prime}}$, where {|$\{\ket{00^{\prime}},\ket{11^{\prime}}\}$ the Schmidt basis of the bi-partite system, and $\alpha\in\mathbb{R}$ a parameter to quantify the degree of entanglement. Alice and Bob perform measurement in the settings $a,b$ and obtain binary outcomes $x,y\in\{0,1\}$.
@@
Figure 1: The experimental setup consists of three observers Alice, Bob, and Maggie ($A,B$, and $M$). The source $S$ emits a pair of entangled quantum particles Alice and Bob perform measurement in the settings $a,b$ and obtain binary outcomes $x,y\in\{0,1\}$. Maggie obtains a variable $z$ with which she can make better predictions about Alice’s outcomes than the standard quantum Born rule.
@@
The third observer possesses a magic prediction box with outcome variables $z$ conditioned on which the predictions of Alice are hypothetically improved. That is, in each run of the experiment, the probabilities for outcomes $a$ given the state $\rho$ may differ from the probabilities for outcomes $a$ given the state $\rho$ plus the value of $z$. The experiment is depicted in Figure 1. In Section 2 a precise assumption will be made to what extent $M$ can share with $A$ the value of $z$.
@@
In the next section the assumptions needed for the theorem are discussed, and the main result presented. It is shown under what conditions quantum mechanics can be considered predictively complete. Signal-locality is shown to suffice to infer empirical completeness for maximally entangled quantum states. However, introducing epistemic boundaries between different users of the same theory—viz. violating the assumption of ‘reliable intersubjectivity’, one can in principle get around limitations on predictive advantage.
Results
@@
Here we formulate the main claim and the assumptions on which it rests. We shall first introduce the auxiliary assumption on the nature of the variables involved to make predictions, i.e. define the variables which are thought of as being ‘accessible’ to observers on which, in turn, predictions supervene. We then make precise what it means for an observer to have predictive advantage, and explicate the necessary assumption on the epistemic relationship between different observers of the theory, which we dub ‘reliable intersubjectivity’. This then allows to make sense of the crucial assumption of signal-locality and the additional constraint of ‘no-conspiracy’ which then serve to derive the main claim of this work. That is, the assumptions cannot all be jointly true.
Formalisation of the assumptions
@@
In the context of our setup there is thus the question what the variables are that Alice, Bob, and Maggie can operationally access, manipulate, and empirically use to condition their physical predictions. By definition we suppose that the set of variables accessible to observers are contained in the manifest domain. This entails the first assumption necessary to formulate the main claim:
@@
This allows to make more precise our second assumption. Consider the setup introduced above (Figure 1) with a bi-partite system with two observers Alice ($A$) and Bob ($B$) and outcomes $x,y$ and measurement settings $a,b$. The prepared quantum state is assumed to be a fixed pure state $\rho$.
@@
The quantum probabilities on the left-hand side depend on Alice’s and Bob’s measurement settings $a,b$, and the quantum state $\rho$. The quantum state leads to a preparation $\mu(z)$ for Maggie’s prediction box. In principle, $z$ can also depend on the settings of the other observers which would lead to a version of superdeterminism, and retro-causation. The no-conspiracy Assumption 5 below rules out such possibilities. $(PA)$ formalises the violation of empirical completeness 1.
The predictive advantage is non-trivial if for at least one value of the variable $z$ coming from the magic prediction box, the prediction deviates from the Born rule for at least some outcome $x$, setting $a$, and state $\rho$, i.e. (omitting $z$’s dependency on other variables)
$\exists z:p(x|a,\rho)=\sum_{y}p(x,y|a,b,\rho)\neq\sum_{y}p(x,y|a,b,\rho,z)=p(x|a,b,\rho,z).$ (2.2)
Maggie is not assumed to perform standard quantum measurements on the shared entangled pair of particles. Thus, the setup of the main claim is different form familiar (extended) Wigner’s friend experiments where it is standardly assumed that some observers (the friends) measure the quantum system and evolve quantum mechanically before super-observers do further quantum operations on their labs.[11, 8, 19] Maggie is hence also not assumed to remain in a superposition of states that is passed on to Alice. We treat the problem as a question about what (informational) relevance the variable $z$ can have for Alice’s outcomes without any assumption what the quantum state, if any, Maggie is in.
Predictive advantage is the assumption that relates her variable to the other variables in the setup. Either Maggie makes the same predictions as Alice—the standard Born probabilities—or, she makes improved predictions in which case $z$ acts as a source for pre-determination of Alice’s outcomes. [^6] Relatedly, in a recent review of Wigner’s friend type arguments Leifer et al. point out that the results show a contradiction between the Born rule (i.e. the frequency by which outcomes will be observed without assumptions about the post measurement state) and unitarity (and not just a tension between unitary evolution and possible state-update rules like the projection postulate)—except for the Frauchiger-Renner argument where the contradiction arises from combining Heisenberg cuts at different levels of observation (see, for example, [27, 38]).
@@
$p(x|a,b,\rho,z)=\sum_{y}p(x,y|a,b,\rho,z).$ (2.3)
The need for reliably intersubjectivity can be framed as reminiscent of the quantum measurement problem.
@@
One natural proposal to (partly) accommodate the apparent ambiguity is to make an assumption about one particular choice of Wigner’s measurement setting (the one with which the original thought experiment is standardly concerned). When he measures in the friend’s basis, i.e. simply asks the friend to report her outcome, then the measurement results are assumed to be compatible. In other words, Wigner essentially ‘checks’ the friend’s reading, and it seems plausible that Wigner’s result equals what the friend observed. Occasionally this assumption is called ‘shared facts’ (e.g. in Adlam’s approach [1]). Moreover, this also follows from the so-called ‘absoluteness of observed events’ $(AOE)$ and ‘observer-independent facts’ (as employed in the extended Wigner’s friend experiments by Brukner et al. [3] and Cavalcanti et al. [8]). [^7] Note that intersubjectivity is distinct from the ‘repeatability’ of quantum measurements. If the same measurement is performed twice, the results are in agreement. Moreover, it also implies that if Wigner performed the same measurement as the friend directly on the target system, their values would come out identical. Compare, for instance, with what Maudlin calls this ‘the problem of effect’ in the context of the quantum measurement problem.[32] The invoked consistency is a feature of the empirical content of the traditional theory and is experimentally well-confirmed.

### v3

# Epistemic Boundaries and Quantum Uncertainty: What Local Observers Can (Not) Predict

@@
Abstract
@@
One of quantum theory’s salient features is its apparent indeterminism, i.e. measurement outcomes are typically probabilistic. We formally define and address whether this uncertainty is unavoidable or whether post-quantum theories can offer a predictive advantage while conforming to the Born rule on average. We present a no-go claim combining three aspects: predictive advantage, no-signalling, and reliable intersubjectivity between quantum observers. The results of the analysis lead to the conclusion that there exists a fundamental limitation on genuine predictive advantage. However, we uncover a fascinating possibility: When the assumption of reliable intersubjectivity between different observers is violated, subjective predictive advantage can, in principle, exist. This, in turn, entails an epistemic boundary between different observers of the same theory. The findings reconcile us to quantum uncertainty as an aspect of limits on Nature’s predictability.
1 Introduction
@@
Insofar as the formalism is deemed valid, quantum theory’s predictions are generally probabilistic. That is, generic physical states imply uncertainty for the outcomes of measurements. At face value, the quantum formalism suggests that its probabilism is unsurmountable, so quantum indeterminism has become one of the characteristics of the theory.
@@
Quantum theory’s indeterminism can be articulated in terms of a ‘decorrelation principle’ [1]: This principle states that when a pure state is assigned to a target system $S$, the outcomes of any measurement are statistically independent of its environment. Consequently, the maximum information that the environment can contain is limited to the information about the target system’s pure quantum state:
$p(x,y|\rho)=p(x|\rho_{S})\cdot p(y|\rho_{E}),$
where $x,y$ are the outcomes of the measurements on system $S$ and the environment $E$, $\rho$ the total joint state, and $\rho_{S},\rho_{E}$ the respective reduced states. [^1] Classical mechanics can be said to obey a similar decorrelation principle. However, the ‘pure’ states are the exact positions and momenta of particles which together with the physical laws contain no uncertainty about the future state of a system. For a caveat to this common claim see [2].
@@
The question arises whether we expect quantum uncertainty to be fundamental in future theories as well, or whether there could exist non-quantum variables with predictive advantage.
@@
But why does the quantum formalism suggest that there is a limit to the predictability of an event? Is it because Nature is fundamentally indeterministic? Still, if Nature were deterministic, this raises the question of why it shouldn’t be predictable. Can quantum uncertainty be a consequence of a more fundamental physical principle?
Apart from the foundational significance, the question has practical relevance and consequences for post-quantum theories: What is the most accurate prediction possible for physical phenomena? What are physical principles independent of the quantum formalism that can be used to make statements about the predictability in post-quantum theories.
@@
In 1935 the puzzle about quantum theory’s completeness culminated in the notorious ‘EPR-argument’: Imagine a widely separated pair of entangled particles that can thus no longer interact with one another. In turn, the situation is said to satisfy Einstein, Podolsky, and Rosen’s element of reality: Since “without in any way disturbing the system” (as the particles are space-like separated) “we can predict with certainty (i.e. with probability equal to unity) the value of a physical quantity, […] there exists an element of physical reality corresponding to this physical quantity.” [3, p. 777]. Since the particles are entangled, one can therefore simultaneously perform a position measurement on one and a momentum measurement on the other particle. If the remote measurement outcome can be predicted with certainty (due to the perfect quantum correlations) without interfering with the system (by assuming the validity of the theory of relativity), the outcome must be pre-determined and thus the quantum description is incomplete: Einstein was “[…] firmly convinced that the essentially statistical character of contemporary quantum theory is solely to be ascribed to the fact that this [theory] operates with an incomplete description of physical systems”. [3, p. 666]
@@
Einstein played with other thought experiments to conceive of scenarios violating Heisenberg’s uncertainty relations (see e.g. perpetua mobilia like the Einstein paradox, [4]). He knew that knowledge of the uncertainty principles doesn’t imply that a more complete theory is impossible. Einstein was primarily concerned with the ontological rather than predictive completeness of quantum mechanics. EPR never claimed that the theory would be predictively complete (see, for instance, [5]).
Important recent results by Barrett, Kent, and Pironio [6] as well as Colbeck and Renner [7] addressed predictive completeness and showed that quantum predictions cannot be extended by employing a version of chained Bell inequalities.
However, these results do not take into account the significance of reliable intersubjectivity, i.e. the shared experience of measurement records by macroscopic observers. Here we demonstrate that reliable intersubjectivity is a crucial and necessary assumption. It can be seen as a generalisation of what is commonly known as ‘absoluteness of observed events’ which plays a crucial role in recent results involving multi-agent scenarios.[8, 9, 10, 11, 12] None of these (extended) Wigner’s friend type experiments [13], however, addresses the question of predictive advantage.
We present a novel argument for the predictive completeness of quantum theory based on no-signalling and reliable intersubjectivity. This is achieved by giving a definition of predictive advantage through a violation of empirical completeness. The definition is purely operational and avoids the introduction of an ontological model [14]. We prove that the quantum state provides complete information about the predictions of measurement outcomes. More concretely, no predictive advantage can objectively exist for all observers of a theory that cannot signal superluminally while reproducing the quantum probabilities on average. However, predictive advantage is indeed possible in the presence of epistemic boundaries: The potential predictive advantage of one observer cannot be shared with other observers.
To our knowledge, this is the first rigorous result on predictive completeness addressing locality and reliable intersubjectivity.
We proceed as follows. Section 2 introduces the relevant notion of empirical completeness and the setup of the main result. In Section 3 we formalise the assumptions of the theorem. The findings are discussed in Section 4, where we comment on related results and the implications for interpretations of quantum theory. The proof of the theorem is given in Appendix A. We conclude in Section 5.
2 Predictions of No-Signalling Observers
@@
Let’s begin with an instructive example from classical physics. Imagine a classical thermodynamic system of a gas in a box to illustrate how empirical incompleteness may turn out. The box of gas may be described as a closed system with a constant temperature. Subject to kinetic gas theory, a measurement of the velocity $v$ of the gas particles in the ensemble is expected to give values distributed according to Maxwell-Boltzmann statistics $f_{MB}(v)$, peaked around some mean velocity. Thus, the velocity $v$ of an individual particle will be found with some likelihood $p(v)$. But let’s imagine the outcomes of many repeated measurements on the same gas yield doubly peaked distributions corresponding to two Maxwell-Boltzmann distributions of different mean velocities. Let’s assume the two peaks originate from the fact that there are two distinct gases in the box, namely nitrogen and oxygen. Since the two gases have different kinetic properties, this leads to separate mean velocities. But to an experimenter without a physical theory distinguishing the two molecules, only one gas is in the box. Thus, the experimenter’s original theory is empirically incomplete since the measurement outcomes will show an in-principle uncertainty resulting from treating the gas as one single substance. A more advanced theory that can distinguish nitrogen from oxygen will obtain more complete predictions. [^2] For this to be possible, an actual experiment, i.e. preparation, has to exist that distinguishes nitrogen from oxygen, such that, on average, the total distribution $f_{MB}(v)$ is still recovered. This is crucial, for otherwise, the improved theory wouldn’t be an empirical completion but rather a metaphysical completion.
@@
All variables involved—outcomes $x$, and preparations $y_{1},y_{2}$—are configurations in the manifest domain $\mathcal{M}$. The set $\mathcal{M}$ is a specification of the directly accessible empirical ‘data’ on which an observer’s knowledge supervenes [^3] Using the standard notion of supervenience, this means that there cannot be a difference in the observers’ knowledge state without a difference in the empirical data. (cf. [1] for details). [^4] It is worth noting that indeed every probabilistic physical theory can be supplemented by a deterministic underpinning. (for any sequence of events there exists a description that contains all information in the initial state of the system, cf. e.g. [1], [15, Chapter 3.4.1]. A viable version of determinism will thus involve restrictions on the domain of variables allowed and the complexity of laws invoked). For instance, quantum theory is explicitly probabilistic, but pilot wave theory provides a deterministic underpinning.
@@
The idea to connect (versions of) locality with quantum completeness dates back to Einstein, Podolsky and Rosen, who argued for the incompleteness of quantum theory based on the existence of entangled states and their non-local properties [3]. If quantum measurement outcomes aren’t further pre-determined by more than the probabilistic quantum prediction, they must be non-local (in the sense of EPR [^5] Let $\rho$ be the state of a space-like separated bi-partite quantum system with measurement results $x$ and measurement settings $a$ on one side, and outcomes $y$ and settings $b$ on the other. Let $\lambda\in\Lambda$ be the supplementing hidden states of the system. Following Shimony [16, 17], two notions of locality are distinguished. Parameter Independence: $p(x|a,b,y,\rho,\lambda)$ $=p(x|a,y,\rho,\lambda),$ (2) $p(y|a,b,x,\rho,\lambda)$ $=p(y|b,x,\rho,\lambda),$ for all $x,y,a,b,\lambda$. That is, the outcome probabilities are independent of the settings of the remote measurement. Outcome Independence: $p(x|a,b,y,\rho,\lambda)$ $=p(x|a,b,\rho,\lambda),$ (3) $p(y|a,b,x,\rho,\lambda)$ $=p(x|a,b,\rho,\lambda),$ for all $x,y,a,b,\lambda$, i.e. the outcome probabilities are independent of the outcomes of the remote measurement. If there exists no hidden state $\lambda$, then outcome independence effectively coincides with EPR’s notion of locality. The conjunction of both parameter and outcome independence is needed to derive Bell’s theorem.) to recover the perfect correlations of entangled states [18]. Supplementing the quantum description with so-called hidden variables could then perhaps restore determinism, thus completing the theory to avoid what is often called ‘action at a distance’.
@@
But by Bell’s theorem, any such completions must be non-local too.[19] To avoid the paradoxical challenges of superluminal signalling this implies that if hidden variables exist, the standard theory can either not be completed, or some mechanism must exist limiting the extent to which such variables are empirically accessible. Otherwise, if experimenters could control hidden variables, they could employ them to signal. A generalisation of this claim is used in the main result of this paper.
@@
The basic setup of the thought experiment used in the main result consists of three observers Alice, Bob, and Maggie which we shall abbreviate with $A,B$, and $M$. The source $S$ emits a pair of spin-$\frac{1}{2}$ particles prepared in the pure entangled state $\rho=\ket{\psi}\bra{\psi}$ | with |$\ket{\psi}:=\alpha\ket{00^{\prime}}+\sqrt{1-\alpha^{2}}\ket{11^{\prime}}$, where {|$\{\ket{00^{\prime}},\ket{11^{\prime}}\}$ the Schmidt basis of the bi-partite system, and $\alpha\in\mathbb{R}$ a parameter to quantify the degree of entanglement. Alice and Bob perform measurements in the settings $a,b$ and obtain binary outcomes $x,y\in\{0,1\}$.
@@
Figure 1: The experimental setup consists of three observers Alice, Bob, and Maggie ($A,B$, and $M$). The source $S$ emits a pair of entangled quantum particles. Alice and Bob perform measurements in the settings $a,b$ and obtain binary outcomes $x,y\in\{0,1\}$. Maggie obtains a variable $z$ with which she can make better predictions about Alice’s outcomes than the quantum Born rule.
@@
The third observer possesses a magic prediction box with outcome variables $z$ conditioned on which the predictions of Alice are hypothetically improved. That is, in each run of the experiment the probabilities for outcomes $x$ given the state $\rho$ may differ from the probabilities for given the state $\rho$ plus the value of $z$. The experiment is depicted in Figure 1. In Section 3 a precise assumption will be made to what extent $M$ can share with $A$ the value of $z$.
@@
In the next section, the assumptions needed for the theorem are discussed, and the main result is presented. It is shown under what conditions quantum mechanics can be considered predictively complete. No-signalling is shown to suffice to infer empirical completeness for maximally entangled quantum states. However, introducing epistemic boundaries between different users of the same theory—viz. violating the assumption of ‘reliable intersubjectivity’, one can in principle get around limitations on predictive advantage.
3 Results
@@
Here we formulate the main claim and the assumptions on which it rests. We shall first introduce the auxiliary assumption on the nature of the variables involved to make predictions, i.e. define the variables which are thought of as being ‘accessible’ to observers on which, in turn, predictions supervene. We then make precise what it means for an observer to have predictive advantage, and explicate the necessary assumption on the epistemic relationship between different observers of the theory, which we dub ‘reliable intersubjectivity’. This then allows us to make sense of the crucial assumption of no-signalling and the additional constraint of ‘no-conspiracy’ which then serve to derive the main claim of this work. That is, the assumptions cannot all be jointly true.
3.1 Formalisation of the Assumptions
@@
In the context of our setup, there is thus the question of what the variables are that Alice, Bob, and Maggie can operationally access, manipulate, and empirically use to condition their physical predictions. By definition, we suppose that the set of variables accessible to observers is contained in the manifest domain. This entails the first assumption necessary to formulate the main claim:
@@
This allows us to make more precise our second assumption. Consider the setup introduced above (Figure 1) with a bi-partite system with two observers Alice ($A$) and Bob ($B$) and outcomes $x,y$ and measurement settings $a,b$. The prepared quantum state is assumed to be a fixed pure state $\rho$.
@@
The quantum state $\rho$ leads to a preparation $\mu(z)$ for Maggie’s prediction box. In principle, $z$ can also depend on the settings of the other observers which would lead to a version of superdeterminism, and retro-causation. The no-conspiracy Assumption 5 below rules out such possibilities. $(PA)$ formalises the violation of empirical completeness 1.
The predictive advantage is non-trivial if for at least one value of the variable $z$ coming from the magic prediction box, the prediction deviates from the Born rule for at least some outcome $x$, setting $a$, and state $\rho$, i.e. (omitting $z$’s dependency on other variables) {dmath} ∃z: p(x|a, b,ρ)=∑_y p(x,y|a,b,ρ) ≠∑_y p(x,y|a,b,ρ,z)=p(x|a,b,ρ, z).
Maggie is not assumed to perform standard quantum measurements on the shared entangled pair of particles. Thus, the setup of the main claim is different from familiar (extended) Wigner’s friend experiments where it is standardly assumed that some observers (the friends) measure the quantum system and evolve quantum mechanically before super-observers do further quantum operations on their labs.[9, 8, 10] Maggie is hence also not assumed to remain in a superposition of states that is passed on to Alice. We treat the problem as a question about what (informational) relevance the variable $z$ can have for Alice’s outcomes without any assumption what the quantum state, if any, Maggie is in.
Predictive advantage is the assumption that relates her variable to the other variables in the setup. Either Maggie makes the same predictions as Alice—the standard Born probabilities—or, she makes improved predictions in which case $z$ acts as a source for pre-determination of Alice’s outcomes. [^8] Relatedly, in a recent review of Wigner’s friend type arguments Leifer et al. point out that the results show a contradiction between the Born rule (i.e. the frequency by which outcomes will be observed without assumptions about the post-measurement state) and unitarity (and not just a tension between unitary evolution and possible state-update rules like the projection postulate)—except for the Frauchiger-Renner argument where the contradiction arises from combining Heisenberg cuts at different levels of observation (see, for example, [20, 21]).
The definition shifts the focus away from the ontological underpinning of the theory—such as e.g. in hidden variable theories—to prediction-making. [^9] Note that there exists a similar but distinct notion of completeness which is concerned with the question of whether the quantum state is ontic. In other words, whether the same state of reality is compatible with more than one wavefunction. (see Pusey, Barrett, and Rudolph’s seminal result on the reality of the quantum state. [22]). We are thus not concerned with the ontological or metaphysical status of the quantum state, but rather with its epistemic role as a means for predicting outcome probabilities.
This supports the need for a clear distinction between what’s manifest, i.e. empirically accessible, from what isn’t, i.e. non-manifest. The variables that may determine a single measurement result don’t necessarily coincide with the ones utilisable for predicting that result. Rather than to attempt restoring determinism in quantum mechanics, the question is whether quantum mechanics is empirically complete.
@@
$p(x|a,b,\rho,z)=\sum_{y}p(x,y|a,b,\rho,z).$ (5)
The need for reliable intersubjectivity can be framed as reminiscent of the quantum measurement problem.
@@
One natural proposal to (partly) accommodate the apparent ambiguity is to assume one particular choice of Wigner’s measurement setting (the one with which the original thought experiment is standardly concerned). When he measures in the friend’s basis, i.e. simply asks the friend to report her outcome, then the measurement results are assumed to be compatible. In other words, Wigner essentially ‘checks’ the friend’s reading, and it seems plausible that Wigner’s result equals what the friend observed. Occasionally this assumption is called ‘shared facts’ (e.g. in Adlam’s approach [23]). Moreover, this also follows from the so-called ‘absoluteness of observed events’ $(AOE)$ and ‘observer-independent facts’ (as employed in the extended Wigner’s friend experiments by Brukner et al. [11] and Cavalcanti et al. [8]). [^10] Note that intersubjectivity is distinct from the ‘repeatability’ of quantum measurements. If the same measurement is performed twice, the results are in agreement. It also implies that if Wigner performed the same measurement as the friend directly on the target system, their values would come out identical. Compare, for instance, with what Maudlin calls ‘the problem of effect’ in the context of the quantum measurement problem.[24] The invoked consistency is a feature of the empirical content of the traditional theory and is experimentally well-confirmed.